A can contains two different kinds of nuts, weighs 5 pounds, and costs $16. One type of nut costs $3.50 a pound. The other type costs $2.75 a pound. How many pounds of $2.75 nuts are there?
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

A can contains two different kinds of nuts, weighs 5 pounds, and costs $16. One type of nut costs $3.50 a pound. The other type costs $2.75 a pound. How many pounds of $2.75 nuts are there?
-------------------
1st nut DATA:
Amount = x lbs ; Value=3.5x dollars
--------------
2nd nut DATA:
Amount = 5-x lbs ; Value=2.75(5-x)=13.75-2.75x dollars
-------------
Mix of nuts DATA;
Amount = 5 lbs ; Value = 16 dollars
----------------
EQUATION:
value + value = 16 dollars
3.5x+13.75-2.75x = 16
0.75x=2.25
x=3 lbs (amount of $3.50 nuts)
5-x=2 lbs ( amount of $2.75 nuts)
----------
Cheers,
Stan H.

Answer by checkley71(8403)

16=3.5x+2.75(5-x)
16=3.5x+13.75-2.75x
16-13.75=.75x
2.25=.75x
x=2.25/.75
x=3 pounds of $3.50 nuts
5-3=2 pounds of $2.75 nuts